# Finite-sample Identification of Continuous-time Parameter-linear Systems

### v1

@@ Abstract
Differentiating noisy, discrete measurements in order to fit an ordinary differential equation can be unreasonably effective. Assuming square-integrable noise and minimal flow regularity, we construct and analyze a finite-difference differentiation filter and a Tikhonov-regularized least squares estimator for the continuous-time parameter-linear system. Combining these contributions in series, we obtain a finite-sample bound on mean absolute error of estimation. As a by-product, we offer a novel analysis of stochastically perturbed Moore-Penrose pseudoinverses.
@@
Errors in variables identification, Filtering and smoothing, Continuous Time System Estimation, Nonlinear system identification, Statistical data analysis
@@
This work has been submitted to IFAC for possible publication.
{pf} Proofs omitted in the submitted version appear like this.
@@ 1 Introduction
This paper presents an estimator with bounded mean absolute error for the parameter vector $\theta$ of the continuous-time parameter-linear system
$\frac{\partial^{m}y}{\partial t^{m}}$ $=\theta^{\intercal}\phi\left(\frac{\partial^{0}y}{\partial t^{0}},\frac{\partial^{1}y}{\partial t^{1}},\ldots,\frac{\partial^{m-1}y}{\partial t^{m-1}}\right)$ (1)
using noisy, discrete measurements of $y$. Our main result, Thm. 15, establishes that $\operatorname{\mathbb{E}}|\hat{\theta}-\theta|$ can be made arbitrarily small by increasing the amount of data collected.
Continuous-time systems arise from physical constitutive relations such as for Newtonian motion $m\ddot{x}=F$, Fickian diffusion $\dot{q}=\Delta q$, capacitance $C\dot{v}=i$, and the like. These principles can be applied without full knowledge of the underlying physical parameters; after defining the dynamic model, one fits the parameters to output trajectories hoping to recover a physical reality. For example, in a capacitor microphone, sound is transduced to capacitance, which is measured in real-time, e.g. by excitation using a high-frequency electrical signal. This interpretation of system-identification-as-measurement applies to magnetic resonance imaging by Maidens et al. (2016), battery diagnostics by Lin et al. (2019), and so on.
Ljung and Wills (2010) rehearse the necessity and difficulty inherent to continuous-time system identification. There is in general no easy translation from continuous to discrete; to discretize (1) by integrating the dynamics accurately would lose the parameter-linear structure. In light of this, the continuous question is not readily amenable to recent answers in finite-sample identification of discrete-time linear (Tsiamis et al., 2022; Ziemann et al., 2023; Matni and Tu, 2019) and parameter-linear (Mania et al., 2022) systems.
A possible solution to problem (1) is to find $\hat{\theta}$ such that the predicted flow generated by $\hat{\theta}^{\intercal}\phi$ best fits the measured $y$. This is the popular prediction error method (PEM) widely endorsed for nonlinear problems (Ljung, 1998, 2010). Some of the PEM writing refers to classical theory on general nonlinear least squares estimation (Chen et al., 2020; Wu, 1981; Lai, 1994; Jennrich, 1969) whose hypotheses are highly technical, and whose conclusions are mostly asymptotic (as opposed to finite).
Prediction error minimization confers no blanket guarantee, and one can devise a toy instance of (1) to make PEM look bad. Take an unforced first-order linear system over a single time step,
$y(0)$ $=y_{0}\neq 0,$
$y(1)$ $=\alpha y(0);$
observed with i.i.d. Gaussian noise:
$Y_{0}$ $=y(0)+\epsilon_{0},$
$Y_{1}$ $=y(1)+\epsilon_{1},$
$\epsilon_{0},\epsilon_{1}$ $\sim\mathcal{N}(0,1).$
Then the maximum-likelihood estimator $\hat{\alpha}=Y_{1}/Y_{0}$ has a Cauchy distribution scaled by $\alpha$. Even with $n\to\infty$ repetitions, there is no approaching the coveted Cramér-Rao bound found in textbooks such as Bard (1974), as the sampling distribution of $\hat{\alpha}_{n}$ does not even have a mean.
On the other hand, we have finite-sample guarantees when system identification amounts to estimating an operator $\Theta$ in
$Y$ $=\Theta X+\text{noise},$ (2)
where generally, $X$ depends on the initial condition, system input, and observable states; $Y$ is an output; and $\Theta$ is an system representation such as a Hankel matrix or transfer function. After multiplying both sides by a generalized inverse of $X$, estimation error in $\Theta$ is bounded in terms of $X$ and the noise profile (Pillonetto et al., 2022; Tsiamis et al., 2022).
What stops us from reducing (1) to (2)? In the latter, the output is linear in the parameters; in the former, the flow is. All we have to do is reconstruct the unmeasured derivatives of $y$; namely, to cast problem (1) into the form
$Y=\Theta\left(X+\text{error}_{X}\right)-\text{error}_{Y},$ (3)
resembling problem (2), we must differentiate the measurements (van Breugel et al., 2020). Differentiation-based identification, according to conventional wisdom, is hindered by bias (due to high-frequency signal) and fluctuation (due to amplification of high-frequency noise). As an overture to the construction of Section 3, we compare it heuristically to PEM on both counts.
First, the derivative filter introduces bias in the form of generalized aliasing: even without noise, it is never possible to get perfect derivatives for a general smooth $y$, because one cannot with certainty impute what happens in between samples. But suppose, as is often the the case, that the sampling period $\Delta t$ is small compared to the regularity of $y$. When applying PEM to such a case, it would seem viable to integrate the flow using a linear multi-step scheme [^1] A good exposition can be found in Dattani (2008). Butcher (2006) is a comprehensive treatment. with steps aligned to the measurement schedule. The quadrature error is controlled by comparing $y$ to local polynomial approximants. Our differentiator will share the local polynomial approximation, but we manage the error systematically in Lemma 6 and beyond.
Second, differentiation amplifies noise. Indeed, the ideal LTI differentiator has Laplace transfer function $s$. But the noisy derivative is a throwaway intermediate representation that gets smoothed in the next step. For example, to estimate $\dot{y}=\theta$, the time average of $\hat{\dot{y}}$, which is the differentiation-based estimate, is very close to the least-squares PEM estimate.
This paper puts the following contributions in series.
  • Analysis of the ridge-regression estimator subject to square integrable noise in both the regressors and predictors. Noise in both variables constitutes an errors-in-variables problem familiar in statistics (Gleser et al., 1987; Hirshberg, 2021; Zhang et al., 2020) and system identification (Barbieri and Diversi, 2021; Khorasani and Weyer, 2021; Fosson et al., 2021). Our stochastic pseudoinverse perturbation bound, Lemma 13, reveals the usefulness of Tikhonov regularization in such problems.
1.1 Notation and conventions
[…]
2 Problem and outline
@@
Our parameter space will be $\Theta=\mathbb{R}^{p}$. The output is $y:[0,T]\times\Theta\to\mathbb{R}$, and the state $\xi$ consists of the derivatives of $y$ of orders 00 through $m-1$. The state dynamics are
@@
with a known initial condition $\xi(0,\theta)=z_{0}$.
@@
The feature map $\phi:\mathbb{R}^{m}\to\mathbb{R}^{p}$ is $\alpha$-Hölder continuous with coefficient $C_{\phi}$.
@@
$\frac{\partial^{m+1}}{\partial t^{m+1}}y$ exists and is Riemann integrable.
It is possible that the system under identification is closed-loop, and in such a case $u$ is embedded in a non-autonomous $\phi$. In such cases, we do not concern ourselves with choosing the best “$u$”: our tools only speak to how good a particular choice is.
2.1 Stochastic observation and estimation
To represent noisy measurement, we use an abstract probability space consisting of outcomes $\Omega$, $\sigma$-field $\mathcal{F}$, and probability measure $\mathbb{P}$. We write $\operatorname{\mathbb{E}}$ for expectation on this probability space.
We have a zero-mean noise process $w:[0,T]\times\Omega\to\mathbb{R}$ such that for each $t\in[0,T]$, $\omega\mapsto w(t,\omega)$ is an independent random variable. There is no need to negotiate with the fine analytical details of continuous-time random processes, as we never do calculus with $w$; however, our bounds require:
Assumption 3
$\operatorname{\mathbb{E}}w(t)^{2}$ is Riemann integrable in $t$.
@@ 2.1 Stochastic observation and estimation
A discrete estimator of $\theta$ is a random variable $\hat{\theta}:\Omega\to\Theta$ in the $\sigma$-field generated by $z_{i,n}$ for all $1\leq i\leq n$.
Definition 5
@@
A sequence of discrete estimators $\{\hat{\theta}_{n}\}_{n\in\mathbb{N}}$ is consistent if $\hat{\theta}_{n}\overset{\mathbb{P}}{\longrightarrow}\theta$.
Our problem is: find a consistent sequence of discrete estimators indexed by $n$, the number of samples.
2.2 Plan
In Section 3 we use finite difference filters to get simultaneous derivatives of $y$ of order 00 through $m$ at almost all of the sample times. The only facts carried forward are the error moments eqq. (15) and (17), which we paraphrase as:
$\sum_{i=1}^{n^{\prime}}\sum_{j=0}^{m}\operatorname{MSE}\widehat{\partial_{t}^{j}y(t_{i})}\lesssim\frac{n^{\prime}}{n^{\frac{2}{2m+3}}},$
where $n^{\prime}$ is the length of filtered output (related to subsampling).
Then in Section 4 we plug these into a $\lambda$-ridge regression estimator to get $\hat{\theta}_{n}$. After a lot of real analysis, we will conclude in Cor. 16 that
$\operatorname{\mathbb{E}}\left\|\hat{\theta}_{n}-\theta\right\|\lesssim n^{-\frac{\alpha}{2m+3}},$
in which the dependence on $\sqrt{n^{\prime}}$ has factored completely through terms of the form $\lambda/\sqrt{n^{\prime}}$.
Thus the paper concludes.
2.3 Open question
We wonder whether our classically fashioned Assumption 3 of square-integrability of $w$ allows for nontrivial strengthening of our main result, either absolutely, in the form of an error moment; or relatively as a tail bound stronger than Mean Absolute Deviation $\operatorname{\mathbb{E}}|\hat{\theta}-\operatorname{\mathbb{E}}\hat{\theta}|$. In particular, does $\hat{\theta}$ exhibit concentration as $n^{\prime}\to n$? What about $n^{\prime}\gg n$ if Section 3 is extended to get inter-sample derivatives?
3 Differentiating noisy signals
Use the $n$ random variables $z_{i,n}$ to fashion $n^{\prime}$ evenly-spaced windows, each containing $N$ consecutive samples. The windows can be chosen to be disjoint, $n^{\prime}=n/N$; overlapping, $n^{\prime}=n-N+1$; or anything in between. Intuitively, a higher $n^{\prime}$ is better (why discard data?), but there is a tradeoff when we get to Theorem 15.
@@ 3 Differentiating noisy signals
Let $n_{i}$ be the first index of the $i$th window. Then the start time of the $i$th window is $t_{i}=n_{i}n^{-1}T$. Let $x^{d}_{i}=\frac{\partial^{d}}{\partial t^{d}}y(t_{i})$, $d=0,1,\ldots,m$. Regard $x_{i}$ as an $\mathbb{R}^{m+1}$ vector, and $\hat{x}_{i}$ likewise (after we construct $\hat{x}_{i}^{d}$). Let $\Delta x_{i}=\hat{x}(t_{i})-x(t_{i})$.
3.1 Building blocks of the filter

### v3

# Estimation Sample Complexity of a Class of Nonlinear Continuous-time Systems

@@ Abstract
We present a method of parameter estimation for large class of nonlinear systems, namely those in which the state consists of output derivatives and the flow is linear in the parameter. The method, which solves for the unknown parameter by directly inverting the dynamics using regularized linear regression, is based on new design and analysis ideas for differentiation filtering and regularized least squares. Combined in series, they yield a novel finite-sample bound on mean absolute error of estimation.
@@
Modelling, Identification and Signal Processing; Machine Learning in modeling, estimation, and control; Estimation
@@ 1 Introduction
Most physical laws are continuous-time. Examples include Newtonian motion $\ddot{x}=m^{-1}F$, Fickian diffusion $\dot{q}=D\Delta q$, and capacitance $\dot{v}=C^{-1}i$. These three relationships are semilinear (in fact, linear) ordinary differential equations in the state variables $x$, $q$, and $v$; with vector fields that are linear in the parameters $m^{-1}$, $D$, and $C^{-1}$, respectively. We view the states as observable and the parameters as unknown. System identification, as a way of measurement, takes on the emphasis of parameter estimation (as opposed to model prediction); this interpretation of system identification has universal application across engineering, e.g. magnetic resonance imaging (Maidens et al., 2016) and battery diagnostics (Lin et al., 2019); and in other areas such as monetary policy econometrics (Chen and Phelan, 2023).
We focus on the case in which all of the state variables are derivatives of the observed output. In chaos-related literature, this structure is called a hyperjerk system. In the geometric control literature, the observed output can also be understood as a flat output of a differentially flat system. In the econometrics and control literature, this structure can be considered a nonlinear continuous-time autoregression with observation noise instead of the more customary process noise.
2 Problem statement
@@
Let $\Theta=\mathbb{R}^{p}$ be parameter space. The output is $y:[0,T]\times\Theta\to\mathbb{R}$, and the state $\xi\in\mathbb{R}^{m}$ consists of the derivatives of $y$ of orders 00 through $m-1$. The state dynamics are
@@
subject to the initial condition $\xi(0,\theta)=\xi_{0}$, where $\phi$ is some known function that does not depend on $\theta$.
@@
The feature map $\phi:\mathbb{R}^{m}\to\mathbb{R}^{p}$ is $\alpha$-Hölder continuous with coefficient $C_{\phi}$. [^1] For a system with exogeneous input, the theory also works with a non-autonomous $\phi$.
@@
$\frac{\partial^{m+1}}{\partial t^{m+1}}y$ exists and is Riemann integrable on $[0,T]$.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space with expectation $\operatorname{\mathbb{E}}$, where $\Omega$ is the sample space, $\sigma$-field $\mathcal{F}$, and probability measure $\mathbb{P}$.
@@
A discrete estimator of $\theta$ is a random variable $\hat{\theta}:\Omega\to\Theta$ expressed as a function of $\{z_{i,n}\}_{1\leq i\leq n}$ that does not depend explicitly on $\theta$.
The goal is to find an estimator design that can be made provably accurate by increasing $n$, which increases the quantity and frequency $T/n$ of data collected.
Problem 5
@@
Find a sequence of estimators $\{\hat{\theta}_{n}\}_{n\in\mathbb{N}}$ such that $\hat{\theta}_{n}\overset{\mathbb{P}}{\longrightarrow}\theta$, with quantitative guarantees for finite $n$.
3 Related work
3.1 Discretization does not solve Problem 5
The discrete-time model $X_{i+1}\approx f(X_{i},\theta)$, where $X$ is state, $f$ is the dynamics, and $\approx$ denotes process or observation noise, is a mainstay in system identification due to its resemblance to nonlinear regression $Y\approx f(X,\theta)$. There is recent progress in finite-sample identification of discrete-time linear (Sarkar et al., 2021; Tsiamis et al., 2022; Ziemann et al., 2023; Matni and Tu, 2019) and parameter-linear (Mania et al., 2022) systems. Can discrete-time identification theory solve the continuous-time Problem 5? The answer is: not easily. An overview of continuous-to-discrete challenges in system identification is available in Ljung and Wills (2010).
One might argue that discretization converts continuous-time models to discrete-time models. But (1) cannot be translated into discrete form without compromising either accuracy or parameter-linearity. For instance, one can obtain the equivalent model
$y((i+1)n^{-1}T,\theta)=f(y(in^{-1}T,\theta),\theta)$ (4)
where the dynamic map $f$ is computed by solving an initial value problem (IVP). But the right side of this equation is no longer linear in $\theta$.
One could take a Taylor expansion of (4) for $n^{-1}T\approx 0$ in order to preserve the parameter-linearity. That would be equivalent to choosing a numerical scheme for integrating the IVP above, with the need to deal with the integration error. This practice would be at least as hard as estimating differentiation error (our approach).
3.2 Curve fitting does not solve Problem 5
In the prediction error method (PEM) for nonlinear system identification (Ljung, 1999; Ljung, 2010; Keesman, 2011, sec. 5.2), we find $\hat{\theta}$ such that the predicted flow generated by $\hat{\theta}^{\intercal}\phi$ best fits the measured $y$. Some of the PEM literature is grounded in classical theory on general nonlinear least squares estimation (Chen et al., 2020; Wu, 1981; Lai, 1994; Jennrich, 1969) In certain large-sample limits, extremum estimators such as PEM and maximum likelihood estimation can be asymptotically consistent.
Theorem 6 (Amemiya [1985, Thm. 4.1.1])
Assume:
  1. 1.
    The parameter space $\Theta$ is a compact subset of $\mathbb{R}^{p}$.
  2. 2.
    $Q_{n}(\omega,\theta)$ is measurable in the outcome $\omega\in\Omega$ and continuous in $\theta\in\Theta$222e.g. a sum-of-squares error between random observation outcome $\omega$ and prediction at $\theta$.
  3. 3.
    $n^{-1}Q_{n}(\omega,\theta)\to Q(\theta)$ in probability uniformly333 i.e. $\lim_{n\to\infty}\operatorname{\mathbb{P}}\left[\sup_{\theta\in\Theta}n^{-1}Q_{n}(\omega,\theta)<\epsilon\right]=1$. in $\theta$ as $n\to\infty$, and $Q(\theta)$ has a unique global maximum at $\theta_{0}$.
Let $\hat{\theta}_{n}(\omega)\in\arg\max_{\theta}Q_{n}(\omega,\theta)$ be a measurable extremum. Then $\hat{\theta}_{n}\overset{\mathbb{P}}{\to}\theta_{0}$.
The global maximum estimator can be difficult to locate (Amemiya, 1985, p. 110), and we give such an example in Section 7 in comparison with our proposed estimator. Furthermore, the global maximum estimator does not come with even asymptotic tail guarantees such as normality. Asymptotic normality requires a number of higher regularity and rank conditions (Amemiya, 1985, Thm. 4.1.3), as well as the following result on local extrema: if $Q_{n}$ has multiple local maxima, one of them converges in $L^{2}(\Omega)$ to a local maxima of $Q$ (Amemiya, 1985, Thm. 4.1.2). On the other hand, our estimator comes with a moment guarantee valid at finite $n$; and it is a “plug-in” estimator that does not require solving any implicit nonlinear equations. Our estimator also works without continuity in $\theta$; see Section 8.
3.3 Linear theory does not solve Problem 5
We have finite-sample guarantees when system identification amounts to estimating an operator $\Theta$ in the linear relationship
$Y$ $=\Theta X+\text{noise},$ (5)
where generally, $X$ depends on the initial condition, system input, and observable states; $Y$ is an output; and $\Theta$ is an system representation such as a Hankel matrix or transfer function. After multiplying both sides by a generalized inverse of $X$, estimation error in $\Theta$ is bounded in terms of $X$ and the noise profile (Ljung, 1999; Sarkar et al., 2021; Pillonetto et al., 2022; Tsiamis et al., 2022; Pintelon and Schoukens, 2012, chap. 11; Bittanti, 2019, sec. 5.5.4.1).
We inherit an impressive legacy of theory and practice on identifying linear systems. Linear system identification, for all practical purposes, has been a solved problem for decades. But the conclusions mostly do not apply to the nonlinear system (1), which does not satisfy scaling or superposition in $y$. That is why we have selected a nonlinear system, (1) which does not satisfy scaling or superposition in $y$, and which enjoys no simplification under taking Laplace transforms.
4 A differentiation-based approach
We attack (1) directly by solving a linear regression for predictors $\frac{\partial^{m}}{\partial t^{m}}y(t)$ in terms of regressors $\phi(\xi(t,\theta))$. This rewrites (1) into the form
$Y=\Theta\left(X+\text{error}_{X}\right)-\text{error}_{Y},$ (6)
differing from (5) only in the errors-in-variables $\text{error}_{X}$.
To get the regressors $\phi(\xi(t,\theta))$ we must differentiate the measurements (Soderstrom et al., 1997; van Breugel et al., 2020). Conventional wisdom holds that differentiation of noisy data is plagued by bias (due to aliasing high-frequency signal) and fluctuation (due to amplification of high-frequency noise). But this impression is fallible on both counts—bias and fluctuation.
Concerning bias: any derivative filter introduces bias in the form of generalized aliasing. Even without noise, it is never possible to get perfect derivatives for a general smooth $y$, because one cannot with certainty impute what happens in between samples. But suppose, as is often the the case, that the sampling period $\Delta t$ is small compared to the regularity of $y$. When applying PEM to such a case, it would seem viable to integrate the flow using a linear multi-step scheme [^4] A good exposition can be found in Dattani (2008). Butcher (2006) is a comprehensive treatment. with steps aligned to the measurement schedule. The quadrature error is controlled by comparing $y$ to local polynomial approximants. Our differentiator will share the local polynomial approximation, but we manage the error systematically in Lemma 9 and beyond.
Concerning fluctuation: any derivative filter amplifies noise, as the ideal LTI differentiator has Laplace transfer function $s$. Our regressor matrix is noisy, but we manage this variability with ridge regularization in Appendix A. [^5] Appendices are only available in the preprint version of this article at Kuang and Lin (2024). To compare with PEM: consider estimation of $\dot{y}=\theta$. The time average of $\hat{\dot{y}}$, which is the differentiation-based estimate, is very close to the least-squares PEM estimate.
4.1 Contributions
After differentiation, our method uses the derivative estimates to invert (1). Our use of finite-difference filters to estimate a continuous-time system resembles Soderstrom et al. (1997), which treats of a linear system with process noise, but our analysis makes the quantitative bias-variance tradeoff explicit. The concept of inverting linear equations to estimate a continuous-time dynamic parameter has a spectral interpretation by Unbehauen and Rao (1997).
A novel analysis of ridge regression subject to square-integrable error in both the regressors and predictors. Noise in both variables is called errors-in-variables in statistics (Gleser et al., 1987; Hirshberg, 2021; Zhang et al., 2020; Pintelon and Schoukens (2012, sec 8.2)) and system identification (Söderström, 2018; Barbieri and Diversi, 2021; Khorasani and Weyer, 2021; Fosson et al., 2021). These problems are ill-behaved because they require the inversion of a random matrix which could a priori be singular with positive probability, leading to heavy tails. Additionally, the phenomenon of regression dilution states that the inverse of a random matrix is too small on average. We account for both of these issues in our analysis.
In brief, let $\sigma>0$ be an unreliable singular value in the data. Ridge regularization picks $\lambda>0$ to control $\left(\sigma+\lambda\right)^{-1}$. The naïve bound is $\left(\sigma+\lambda\right)^{-1}\leq\lambda^{-1}$ which is an $L^{\infty}$-type bound that is forced by the worst-case scenario $\sigma=0$. Rather, we use a very detailed $L^{1}$-type pseudoinverse perturbation bound, Lemma 16 (Appendix A). This is the main idea in the proof of Thm 8.
5 Preliminaries
[…]
6 Solution to Problem 5
This section constructs an estimator that solves Problem 5.
6.1 Differentiation
Partition the $n$ random variables $z_{j}$ into to $n^{\prime}$ windows, each containing $N$ consecutive samples, where $N$ will be determined later. The windows can be chosen to be disjoint, $n^{\prime}=n/N$; overlapping by $n-1$ points, $n^{\prime}=n-N+1$; or anything in between.  [^6] Sharpening the analysis to cover overlapping-window designs is an area of ongoing work.
@@ 6.1 Differentiation
Let $n_{i}$ be the first index of the $i$th window. Then the start time of the $i$th window is $t_{i}=n_{i}n^{-1}T$. Augment $\xi$ of (1) with an extra derivative to define $x^{d}_{i}=\frac{\partial^{d}}{\partial t^{d}}y(t_{i})$, $d=0,1,\ldots,m$. Regard $x_{i}$ as an $\mathbb{R}^{m+1}$ vector.
Theorem 7
For some choice of $N$, there exist statistics $\hat{x}_{i}^{d}$ for all $i\in[1\ldots n^{\prime}]$ and $d\in[0\ldots m]$, given as linear functions of the data $z_{j}$, $j\in[1\ldots n]$, that satisfy
$\operatorname{\mathbb{E}}\sum_{i=1}^{n^{\prime}}\sum_{d=0}^{m}\left\|\hat{x}_{i}^{d}-x_{i}^{d}\right\|^{2}$ $\leq\epsilon n^{\prime}n^{-\frac{2}{2m+3}},$
where $\epsilon>0$ is independent of $n$.
This counterintuitive result on tuning arbitary-order finite-difference filters for noisy data may be the first of its kind suggests that noisy data can be practically differentiated for statistical use in a certain large-sample limit.
The proof using orthogonal polynomials is constructive and given in Appendix A. Henceforth the notation $\hat{x}_{i}^{d}$ will refer to the estimators provided by Theorem 7.
6.2 Regularized least squares
Let $P$ be the projection matrix onto the first $m$ entries. Introduce the shorthand $\phi_{i}=\phi(Px_{i})$ and $\hat{\phi}_{i}=\phi(P\hat{x}_{i})$.
Let $\Phi\in\mathbb{R}^{n^{\prime}\times p}$ be the matrix having $\phi_{i}^{\intercal},i=1,2,\ldots,n^{\prime}$ as rows, and let $\hat{\Phi}$ be the same, with $\hat{\phi}_{i}$ replacing $\phi_{i}$. Let $\lambda\in(0,\sigma_{\text{min}}(\Phi))$ be a regularization coefficient to be determined later. Define $R,\hat{R}\in\mathbb{R}^{(p+n^{\prime})\times p}$ as the block matrices
Define $u,\hat{u}\in\mathbb{R}^{p+n^{\prime}}$ as
$u$ $=\begin{pmatrix}0_{p\times 1}\\ x^{m}_{1}\\ x^{m}_{2}\\ \vdots\\ x^{m}_{n^{\prime}}\end{pmatrix},\quad\hat{u}=\begin{pmatrix}0_{p\times 1}\\ \hat{x}^{m}_{1}\\ \hat{x}^{m}_{2}\\ \vdots\\ \hat{x}^{m}_{n^{\prime}}\end{pmatrix}.$ (8)
$\theta$ $=R^{\dagger}u$
which motivates the estimator
$\hat{\theta}$ $=\hat{R}^{\dagger}\hat{u}.$
Theorem 8
For some choice $\lambda=\lambda(n)$, $\operatorname{\mathbb{E}}\left\|\hat{\theta}-\theta\right\|=O(n^{-\alpha/(2m+3)})$ as $n\to\infty$.
This result is the first finite-sample frequentist guarantee for parameter estimation of a large class of nonlinear continuous-time systems from discrete data.
The proof is constructive and given in Appendix B.
7 Numerical example
Consider the problem of estimating the coefficients $\theta=(\theta_{1},\theta_{2})$ in the Van der Pol oscillator:
$\ddot{x}(t)$ $=\theta_{1}(1-x(t)^{2})\dot{x}(t)+\theta_{2}x(t),\quad 0\leq t\leq T$
The measurements are i.i.d. Normal random variables
$z_{i}$ $\sim\mathcal{N}(x(in^{-1}T),\sigma^{2}).$
Constants pertaining to the problem and the estimator are available in Table 3.
For each of four different values of $N$, we constructed an estimator $\hat{\theta}$ using the filtered least squares idea (with disjoint windows) presented in this paper and used 10000 trials to estimate its sampling distribution. The marginal distributions for $\theta_{1}$ and $\theta_{2}$ are presented in Figures 1 and 2, respectively.
We see that a higher value of $N$, which corresponds to more aggressive smoothing, decreases the estimator’s variance while increasing its bias. The $L^{2}$ error is also minimized at $N=200$, as we can see from absolute (Table 1) and relative errors (Table 2).
In order to illustrate why a direct method such as our $\hat{\theta}$ is apt for this problem, we point to the fact that the negative least squares objective, which corresponds to maximum likelihood, is two-dimensional and far from convex, which renders PEM practically unusable in this case. Using the noisy data from one trial, we computed the likelihood function $\mathcal{L}(\theta)$ and plotted it with $\theta_{2}$ and $\theta_{1}$ held constant at their true values in Figures 3 and 4, respectively. These figures show that while the likelihood function attains a maximum at the true parameter value, the likelihood landscape is rife with local maxima, and maximum likelihood estimation is unreliable without strong prior information.
It is a promising two-step estimation scheme to compute $\hat{\theta}$ via our plug-in estimator, then use the rough estimate to initialize a local search algorithm for the maximum likelihood or maximum a posteriori point estimate.
Figure 1: Sampling distribution of $\hat{\theta}_{1}$ at different choices of $N$.
Figure 2: Sampling distribution of $\hat{\theta}_{2}$ at different choices of $N$.
Figure 3: Log-likelihood of $\theta_{1}$ at the true $\theta_{2}$.
Figure 4: Log-likelihood of $\theta_{2}$ at the true $\theta_{1}$.
$N$ $\text{RMSE}(\hat{\theta}_{1})$ $\text{RMSE}(\hat{\theta}_{2})$ $\text{RMSE}((\hat{\theta}_{1},\hat{\theta}_{2}))$
50 3.77 72.74 5305.53
100 1.00 18.45 341.26
200 0.88 6.37 41.37
400 3.82 13.69 202.03
Table 1: Experimentally determined root mean squared error (RMSE) of $\hat{\theta}$ at different values of $N$.
$N$$\text{\%RRMSE}(\hat{\theta}_{1})$$\text{\%RRMSE}(\hat{\theta}_{2})$$\text{\%RRMSE}((\hat{\theta}_{1},\hat{\theta}_{2}))$
50 9.43 18.19 18.12
100 2.50 4.61 4.60
200 2.20 1.59 1.60
400 9.54 3.42 3.54
Table 2: Relative root mean squared error (RMSE) of $\hat{\theta}$ at different values of $N$, obtained by normalizing columns of Table 1 by $|\theta_{1}|$, $|\theta_{2}|$, and $\left\|\theta\right\|$, respectively.
Constant Meaning Value
$\theta_{1}$ parameter (nonlinear damping) 40
$\theta_{2}$ parameter (squared frequency) -400
$T$ end time 1
$x(0)$ initial displacement 1
$\dot{x}(0)$ initial rate 20
$\sigma^{2}$ observation variance $10^{-4}$
$n$ number of samples 10000
$N$ window size 50, 100, 200, 400
$\lambda$ regularization coefficient 1
Table 3: Values of constants in simulating a Van der Pol oscillator and estimating its parameter.
8 Theoretical example
Let the system model be any instantiation of (1) with $\phi$ satisfying $\phi(0)=0$, $\nabla\phi(0)\neq 0$:
Let $t_{i}$ be the observation times. Let the observations obey
$z_{i}$ $=y(t_{i})+w_{i},$ (10)
where the $w_{i}$ are independently distributed as
$w_{i}$ =$=\begin{cases}e+y(t_{i})&\text{w.p. }1/2\\ -e-y(t_{i})&\text{w.p. }1/2\end{cases}$ (11)
for some $e\in\mathbb{R}$. Because the noise $w_{i}$ is an atomic random variable that depends on $\theta$, a likelihood function does not exist; any solution has to rely on smoothing. Our method, with its accuracy guarantees, still applies, because $\operatorname{\mathbb{E}}\left\|w_{i}\right\|^{2}=\left\|e+y(t_{i})\right\|^{2}<\infty$.
This example also shows the need for ridge regularization. For all sufficiently small $\epsilon$, with probability $2^{-n}$, every row in the regressor matrix $\hat{\Phi}$ will be equal to the same small vector of magnitude $|e|$. Accordingly, the unregularized pseudoinverse of $\hat{\Phi}$ will have a singular value on the order of $|e|^{-1}$, which can cause $\hat{\theta}$ to blow up. Such an event, in which $\theta$ is unidentifiable, is exponentially unlikely. Ridge regularization serves as a kind of insurance that truncates this tail risk, so that the estimation error is at least finite.
9 Conclusion
Continuous-time parameter-linear systems are widely useful, but they do not have an estimation theory with provable finite-sample accuracy. While smoothed least squares is a mainstay in linear system identification, we have contributed a theoretical basis for optimal smoothing and regularization in a nonlinear problem in which traditional proof techniques centered on transfer function algebra do not apply. By using the parameter-linear structure to reformulate the parameter estimation problem as linear least squares, we are able to use a stochastic continuity argument to prove $L^{1}$ consistency in the limit as $n\to\infty$ with $T$ held constant. This approach even works for problems with discontinuous observation noise in which maximum-likelihood large-sample theory does not apply.
Theorem 8 is likely the first identification result in the literature with
  1. 1.
    quantitative, finite-sample consistency, for
  2. 2.
    a large class of nonlinear systems (1) with continuous and unbounded parameter,
  3. 3.
    observed at discrete times, with
  4. 4.
    heavy-tailed noise.
The mathematical technology underlying this result is interesting in its own right, and contains new methods for analyzing differentiation filters and regularized least squares.
Appendices
Appendices A through C are available in Kuang and Lin (2024), the preprint version of this paper, available: https://arxiv.org/abs/2312.05382.
Appendix A Proof of Theorem 7
A.1 Building blocks of the filter